OpenStudy (anonymous):

a school wants to build a rectangular playground that will have a diagonal length of 75 yards. how wide can the playground be if the length has to be 30 yards?

OpenStudy (anonymous):

Isn't it 75-30??

OpenStudy (anonymous):

no it uses the pythagorean theorem.

OpenStudy (jdoe0001):

 \(\Large c^2=a^2+b^2\implies \sqrt{c^2-a^2}=b\)

OpenStudy (anonymous):

yes that how it is.

OpenStudy (anonymous):

i got the answer of 68.7 but after i rounded to nearest tenth.@jdoe0001 is it right?

DivineSolar (divinesolar):

Book 15 show me what you plugged in and I will solve for you as well to see if your answer is correct. as in the full equation sqrt72?^2

OpenStudy (anonymous):

okay so 75 is the diagonal length. and the measure of one of the legs is 30. so i used the a^2 + B^ = c^2 so 30^ + b^2 = 70^ , 900+ B^2= 5,625 so i subtract 5625-900= 4725. then i square 4,725 which equals 68.7 rounded to nearest tenth.

OpenStudy (anonymous):

so thats what i did to do it. @DivineSolar

DivineSolar (divinesolar):

68.7 when you square 4.725 is correct.

DivineSolar (divinesolar):

Also you meant 30^2? Right?

OpenStudy (anonymous):

yes

DivineSolar (divinesolar):

mkay one moment.

DivineSolar (divinesolar):

Your answer is wrong.

DivineSolar (divinesolar):

Lets try this again okay step by step?

OpenStudy (anonymous):

yes

DivineSolar (divinesolar):

Hold on. I may have made an error? redoing the equation.

OpenStudy (anonymous):

oops did u count the 75 as a hypotenuse. because the first time i type the equation i did it fast.

DivineSolar (divinesolar):

Okay I will give you all the breathtaking steps.

DivineSolar (divinesolar):

I love my application, here is a screenshot.

DivineSolar (divinesolar):

https://gyazo.com/a845675dafceabc3153d93836442b1af

DivineSolar (divinesolar):

This is based off your equation.

OpenStudy (anonymous):

u mean 30^2+b^2=75^2

DivineSolar (divinesolar):

Scroll back up. To look and see the equation is c^2 = a^2 + b^2

OpenStudy (anonymous):

yeah it right but i looked at my paper and the hypotenuse is actually 75.

DivineSolar (divinesolar):

With your equation we do get the final answer.

DivineSolar (divinesolar):

as 68.7 since we get \[b = -15 \sqrt{21}\]

DivineSolar (divinesolar):

However with c^2 = a^2 + b^2 we get \[b = -20 \sqrt{10}\]

DivineSolar (divinesolar):

And with that we get the decimal b=−63.245553

OpenStudy (anonymous):

so was i wrong?

DivineSolar (divinesolar):

Do you have any answer choices by any chance?

DivineSolar (divinesolar):

Or is this specifically a word problem?

OpenStudy (anonymous):

no u just have to solve it.

DivineSolar (divinesolar):

Okay in your class do we use your formula or my formula, if its yours then yes your final answer is correct, if it is my formula then my answer is the correct one ^_^